# Online Resource Allocation via Static Bundle Pricing

### v1

@@ 2 Preliminaries
We use the notation $x]\coloneqq\{1,\ldots,x\}$ for any $x\in\mathbb{Z}_{>0}$. For $v\in\mathcal{S}$ define $q_{b,v}\coloneqq\mathbb{P}_{v_{b}\sim\mathcal{D}_{b}}\!\left[v_{b}=v\right]$ and let $\mathcal{D}\coloneq\prod_{b\in[n]}\mathcal{D}_{b}=\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$. Let $I\sim\mathcal{D}$ be an instance of buyers’ values. For each buyer $b$, we denote by $v_{b}$ the random variable corresponding to its value. Note that, throughout the paper, whenever an index of the variables $x_{b,v}$ where $b\in[n],\ v\in\mathcal{S}$ is omitted, we implicitly assume summation over that index (unless stated otherwise). For example, $x_{b}=\sum_{v\in\mathcal{S}}x_{b,v}$ and $x_{v}=\sum_{b\in[n]}x_{b,v}$. For an event $\mathcal{E}$ we define $\mathbbm{1}{\left\{\mathcal{E}\right\}}$ as its indicator function. Finally, let $\ell_{A}(e)$ denote the load of item $e$ due to buyers in $A$, i.e. $\ell_{A}(e)=\sum_{b\in A}\mathbbm{1}{\left\{e\in S_{b}\right\}}$.
2.1 Linear Program
Like in many other works in the stochastic setting, we will compare against the expected offline optimal social welfare Opt through a fractional relaxation of the problem. Consider the following LP relaxation:
$\max\quad$ $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}x_{b,v}\cdot v$
s.t. $\sum_{b\in[n]:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}\leq c(e),\qquad\forall e\in\mathcal{M}$
$0\leq x_{b,v}\leq q_{b,v},\qquad\forall b\in[n],\ \forall v\in\mathcal{S}$
@@ 2.1 Linear Program
Note that in the above LP, the capacity constraints are enforced only in expectation (ex-ante), so $x$ represents a feasible ex-ante fractional allocation. Let FracOpt be the optimal value of the previous LP. Generally, fractional relaxations upper bound the optimal welfare Opt, (see for instance Lemma 2.1 in [chawla18]). For completeness, in Appendix G we prove the following.
Lemma 2.1.
$\textnormal{{FracOpt}}\geq\textnormal{{Opt}}$.
2.2 Structure of the LP solution
Before analyzing our bundle pricing, we first establish several structural properties of the fractional optimal solution of the LP. The key conceptual tool for understanding this structure is the notion of value exchange.
Suppose for the moment that we ignore the constraints $x_{b,v}\leq q_{b,v}$. Consider two variables $x_{b^{\prime},v^{\prime}}>0$ and $x_{b^{\prime\prime},v^{\prime\prime}}$ with $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$. If we decrease $x_{b^{\prime},v^{\prime}}$ by a small amount (mass) $\delta>0$ and increase $x_{b^{\prime\prime},v^{\prime\prime}}$ by the same amount, then the capacity constraints remain satisfied, since allocating mass on a subset of a set consumes no more capacity than allocating the same mass on the original set. This exchange strictly increases the LP objective whenever $v^{\prime\prime}>v^{\prime}$, as we shift mass from a lower-value variable to a higher-value one without affecting feasibility. Thus, in an optimal LP solution $x$, no such improving exchange should be possible. The only way such an exchange can be blocked is by the constraints $x_{b^{\prime\prime},v^{\prime\prime}}\leq q_{b^{\prime\prime},v^{\prime\prime}}$. Consequently, for two variables $0<x_{b^{\prime},v^{\prime}}$ and $x_{b^{\prime\prime},v^{\prime\prime}}$ with $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$, whenever $v^{\prime\prime}>v^{\prime}$, we must have $x_{b^{\prime\prime},v^{\prime\prime}}=q_{b^{\prime\prime},v^{\prime\prime}}$. In the following section, we formalize this intuition, first for each buyer individually and then for each set $S$.
Per Buyer.
Intuitively, our goal is to determine a price to offer each buyer. If we did not require the mechanism to be anonymous or static, a natural approach would be the following: upon the arrival of buyer $b$ with realized value $v$, we accept them with probability $x_{b,v}/q_{b,v}$. Recall that $x_{b,v}$ can be interpreted as the ex ante probability mass of allocating $S_{b}$ to buyer $b$ when $v_{b}=v$. Thus this acceptance rule ensures that buyer $b$ is served with unconditional probability $q_{b,v}\cdot(x_{b,v}/q_{b,v})=x_{b,v}$, matching the LP’s marginal allocation in expectation. Of course, this does not yield a feasible mechanism, because the LP capacity constraints hold only in expectation and may be violated ex-post. The subsequent analysis will show how to overcome this limitation.
From this perspective, the LP solution guides our pricing decisions. Interpreting acceptance probabilities as price thresholds, the LP suggests which values should be priced out and which should be encouraged. If the LP assigns zero weight to a particular value, we should not accept buyers at that level, suggesting a higher price in a static-price mechanism. Conversely, if the LP assigns the maximum feasible weight to a value, this suggests that the corresponding price should not be set above that level. However, there may exist some intermediate crucial values where neither condition applies. Properly handling these crucial values is essential to the design of our mechanism.
@@ 2.2 Structure of the LP solution
Formally, fix a buyer $b$ and consider the vector $X^{(b)}=[x_{b,v}]_{v\in\mathcal{S}}$, whose entries correspond to the optimal fractional LP allocation for buyer $b$, ordered in increasing order of $v$.
Definition 2.2.
A value $v\in\mathcal{S}$ is tight for buyer $b$ if $x_{b,v}=q_{b,v}>0$, and crucial if $0<x_{b,v}<q_{b,v}$.
The following lemma, proved in Appendix G, characterizes the structure of the vector $X^{(b)}$. The proof relies on the notion of value exchange introduced above.
@@
  • a (possibly empty) prefix of values $v$ with $x_{b,v}=0$,
  • followed by at most one value $v$ with $0<x_{b,v}<q_{b,v}$ (the crucial value),
  • followed by a (possibly empty) suffix of values $v$ with $x_{b,v}=q_{b,v}$ (the tight values).
Per Set S.
A central challenge in our mechanism is to offer identical price menus to all buyers. While identical pricing is crucial for maintaining anonymity, it must be combined with additional structural properties, such as subadditivity across sets, which we address in later sections. In this subsection, we focus on the implications of identical pricing. Specifically, anonymity allows us to treat all buyers interested in the same set $S$ uniformly under the mechanism.
For a set $S$, let $\mathcal{A}_{S}\coloneq\{b:S_{b}=S\}$ be the set of buyers that desire $S$. For $v\in\mathcal{S}$ let $x_{S,v}=\sum_{b\in\mathcal{A}_{S}}x_{b,v}$ (define $q_{S,v}$ similarly). Denote $X^{(S)}=[x_{S,v}]_{v\in\mathcal{S}}$ (where entries are ordered in increasing order of $v$ as before). Lastly, define $\textnormal{{FracOpt}}_{S}=\sum_{v\in\mathcal{S}}x_{S,v}\cdot v$. In Appendix G we prove the following, with crucial and tight values of $S$ defined similarly to the single buyer case.
@@
  • a (possibly empty) prefix of values $v$ with $x_{S,v}=0$,
  • followed by at most one value $v$ with $0<x_{S,v}<q_{S,v}$ (the crucial value),
  • followed by a (possibly empty) suffix of values $v$ with $x_{S,v}=q_{S,v}$ (the tight values).
We next introduce a definition that will be essential for setting our posted prices.
Definition 2.5.
The largest value $w\in\mathcal{S}$ such that $x_{S,w}<q_{S,w}$ is called the important value of $S$ and is denoted by $w_{S}$. Note that such value always exists since we can enforce $x_{S,0}=0$. Observe also that if $S$ has a crucial value, then its important value is equal to the crucial value of $S$.
3 General Algorithmic Framework
We begin by introducing a general algorithmic framework. This framework serves as the foundation for all subsequent sections, where we instantiate it in several settings: the general single-minded case, single-minded case with maximum bundle size at most $d$ and general graphs.
3.1 A High-Level Algorithm
We present a two-step framework that converts the scaled LP solution into a static and anonymous (possibly randomized) bundle pricing. We first scale down the capacities, and then translate the per-set structure of the scaled solution into a menu of posted prices for each desired set $S$ (note that a set $S$ is desired if there exists a buyer $b$ such that $S_{b}=S$). At a high level, our framework consists of two main components.
Step 1: Scale down the LP.
We first solve the original LP but with the capacity constraints scaled down by a factor of $\gamma>1$ depending on the setting, i.e. $c(e)$ in the first LP constraint becomes $c(e)/\gamma$. Denote by $\textnormal{{FracOpt}}_{\gamma}$ the new fractional optimum. Note that $\textnormal{{FracOpt}}_{\gamma}\geq\textnormal{{FracOpt}}/\gamma$
Step 2: Post the menu M.
We now create the randomized menu $M$ from the solution of the fractional relaxation according to Algorithm˜1. Our menu is randomized in the sense that the algorithm draws a menu from a distribution before arrivals; importantly, the menu itself does not contain lotteries. The construction is guided by the per-set structure of the scaled LP. For values with $x_{S,v}=0$, we do not offer $S$ at a price that would attract these buyers. For tight values with $x_{S,v}=q_{S,v}$, we offer $|\mathcal{A}_{S}|$ copies so that we effectively accept all such buyers when they appear, with item availability (rather than menu-entry multiplicity) being the only potential bottleneck. Finally, when $S$ has a crucial value, we use a more careful (and in some cases randomized) offer at price $w$ to accept these buyers in expectation close to the mass prescribed by the LP. We are now ready to define our algorithm.
@@ 3.1 A High-Level Algorithm
1:Optimal solution $x$ of the scaled-down LP
2:for each (desired) bundle $S$ do
3:  Let $w$ be the important value of $S$ (w.r.t. to $x$).
4:  Offer $|\mathcal{A}_{S}|$ copies of $S$ at price $w+1$.
5:  if $S$ has crucial value $w$ and $x_{S,w}\cdot w>\displaystyle\sum_{v\in\mathcal{S}:v>w}x_{S,v}\cdot v$ then
6:   if $x_{S,w}>1$ then
7:     Offer $\lfloor x_{S,w}\rfloor$ additional copies of $S$ at price $w$.
8:   else
9:     With probability $\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$, offer one additional copy of $S$ at price $w$.
10:   end if
11:  end if
12:end for
Definition 3.1 (Algorithm).
Fix a menu $M$ and an instance of buyers’ realized values $I$. Our algorithm, denoted $\textnormal{{Alg}}_{M,I}$, operates as follows.
@@
    The adversary selects an arrival order of the buyers in $\mathcal{N}$.
@@
    Buyers arrive sequentially. When buyer $b$ arrives, they examine the current menu $M$ and choose a subset $J$ of available menu entries that maximizes their utility (possibly empty), subject to feasibility with the remaining item copies.
  3. 3.
    If $J\neq\emptyset$, then for each chosen entry $(T_{j},p_{j},c_{j})$ we decrease its remaining multiplicity by $1$, and for each item $e\in T_{j}$ we decrease the remaining copies of $e$ by $1$. Buyer $b$ is deemed accepted.
We will use $\textnormal{{Alg}}_{M,I}$ to denote either the set of buyers accepted by our algorithm or the total value they contribute; the intended meaning will be clear from the context. Our goal is to prove that $\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\geq\alpha\cdot\textnormal{{FracOpt}}$ for some $\alpha>0$.
Menu Subadditivity.
Throughout all subsequent sections, we will assume that buyers in $\mathcal{A}_{S}$ purchase only menu entries that sell exactly the bundle $S$. This is without loss of generality: in Appendix B we show that our menu prices are essentially subadditive, hence no buyer who desires $S$ can increase their utility by purchasing entries corresponding to bundles different from $S$.
Tie-Breaking.
We will also assume that buyers break ties in favour of purchasing, thus if they have value $v$ they will purchase a copy at price $v$. In Appendix B we remove this assumption with a small price perturbation.
3.2 The Unconstrained Algorithm
We now introduce an analytical tool, the Unconstrained Algorithm. This is a hypothetical version of our mechanism that completely ignores item capacities and only enforces the constraints imposed by the menu. In particular, the menu may induce total demand that exceeds the available copies of some items, and the Unconstrained Algorithm still serves all such requests. We use this algorithm purely for analysis: it is not required to be feasible, but it allows us to bound the performance of our actual (capacity-feasible) mechanism by comparing their expected welfare. More formally:
Definition 3.2 (Unconstrained Algorithm).
Fix a menu $M$ and an instance of buyers’ realized values $I$. Our unconstrained algorithm, denoted $\textnormal{{UAlg}}_{M,I}$, operates as follows.
  1. 1.
    Buyers arrive sequentially according to the order that the almighty adversary selected for $\textnormal{{Alg}}_{M,I}$. When buyer $b$ arrives, they examine the current menu $M$ and choose a subset $J$ of available menu entries that maximizes their utility (possibly empty). Item capacities are ignored.
  2. 2.
    If $J\neq\emptyset$, then for each chosen entry $(T_{j},p_{j},c_{j})$ we decrease its remaining multiplicity by $1$. Buyer $b$ is deemed accepted.
We will use $\textnormal{{UAlg}}_{M,I}$ to denote either the set of buyers accepted by the unconstrained algorithm or the total value they contribute; the intended meaning will be clear from the context. Note that a buyer that belongs to $\textnormal{{Alg}}_{M,I}$ must belong to $\textnormal{{UAlg}}_{M,I}$. We next state a key guarantee for UAlg, proved in Appendix C, showing that the unconstrained algorithm captures a constant fraction of $\textnormal{{FracOpt}}_{\gamma}$.

### v2

@@ 4 Notation and Preliminaries
We let $\mathcal{D}\coloneq\prod_{b\in[n]}\mathcal{D}_{b}=\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$ and let $I\sim\mathcal{D}$ be an instance of buyers’ values. For each buyer $b$, we denote by $v_{b}$ the random variable corresponding to its value. For a set $S$, let $\mathcal{N}_{S}\coloneq\{b:S_{b}=S\}$ be the set of buyers that desire $S$. We let $q_{S,v}=\sum_{b\in\mathcal{N}_{S}}q_{b,v}$ be the total mass over buyers in $\mathcal{N}_{S}$ that have realized value $v$. Whenever an index of the variables $x_{b,v}$ is omitted, we implicitly assume summation over that index (unless stated otherwise). E.g., $x_{b}=\sum_{v\in\mathcal{U}}x_{b,v}$ and $x_{v}=\sum_{b\in[n]}x_{b,v}$. For an event $\mathcal{E}$, we let $\mathbbm{1}{\left\{\mathcal{E}\right\}}$ denote its indicator function. We sometimes use $\ell_{A}(e)$ to denote the load of item $e$ due to buyers in set $A\subseteq\mathcal{N}$, i.e. $\ell_{A}(e)=\sum_{b\in A}\mathbbm{1}{\left\{e\in S_{b}\right\}}$.
5 Ex-Ante Linear Program and Solution Structure
The competitive ratio of our mechanism is established against (and our pricing menu is guided by) an upper bound on the expected optimal social welfare Opt given by the following ex-ante fractional relaxation of the social welfare maximization problem:
$\max\quad$ $\sum_{b\in[n]}\sum_{v\in\mathcal{U}}x_{b,v}\cdot v$
s.t. $\sum_{b\in[n]:e\in S_{b}}\sum_{v\in\mathcal{U}}x_{b,v}\leq c(e)\qquad\forall e\in\mathcal{M}$
$0\leq x_{b,v}\leq q_{b,v}\qquad\forall b\in[n],\ \forall v\in\mathcal{U}$
@@ 5 Ex-Ante Linear Program and Solution Structure
In the above linear program, usually referred to as ex-ante LP (or EA-LP, in short), the capacity constraints are enforced only in expectation (ex-ante), so its solution represents a feasible ex-ante fractional allocation. We let FracOpt denote EA-LP’s optimal value. Generally, ex-ante fractional relaxations upper bound the prophet’s optimal expected social welfare Opt (see e.g., (Chawla et al., 2019, Lemma 2.1)). For completeness, in Appendix F, Lemma F.1, we prove that $\textnormal{{FracOpt}}\geq\textnormal{{Opt}}$.
5.1 Structure of Optimal Fractional Allocation
Before formally describing our bundle pricing scheme, we provide intuition and establish some structural properties of the optimal EA-LP allocation, denoted $(x_{b,v})_{b\in[n],v\in\mathcal{U}}$ in what follows.
If we did not require the mechanism to be anonymous or static, a natural approach would be to accept each new buyer $b$ with realized value $v$ with probability $x_{b,v}/q_{b,v}$. Namely, $x_{b,v}$ is interpreted as the ex-ante probability of allocating $S_{b}$ to buyer $b$, if $b$’s realized value is $v_{b}=v$. This allocation ensures that buyer $b$ is accepted with probability $x_{b}=\sum_{v\in\mathcal{U}}x_{b,v}=\sum_{v\in\mathcal{U}}q_{b,v}\cdot(x_{b,v}/q_{b,v})$ in total, matching the allocation suggested by EA-LP. However, this allocation is not anonymous (because buyers $b$ and $b^{\prime}$ that demand the same bundle $S$ and have the same realized values $v$ may be accepted with different probability due to different priors $q_{b,v}$ and $q_{b^{\prime},v}$). Moreover, such an allocation may not be feasible, because the item capacity constraints may be violated ex-post. In the following, we show how to implement a static and anonymous pricing scheme which mimics this allocation.
Fractional Allocation per Buyer.
Following up on the intuition above, we call a value $v\in\mathcal{U}$ tight for a buyer $b$ if $x_{b,v}=q_{b,v}>0$, and crucial, if $0<x_{b,v}<q_{b,v}$. Intuitively, we should always (resp. sometimes) accept $b$ if their value is tight (resp. crucial).
@@ 5.1 Structure of Optimal Fractional Allocation
For each buyer $b$, we consider the vector $X^{(b)}=(x_{b,v})_{v\in\mathcal{U}}$, whose entries correspond to the optimal fractional allocation to buyer $b$, ordered in increasing order of $v$. We have that:
@@
  • A (possibly empty) prefix of values $v$ where $x_{b,v}=0$,
  • followed by at most one value $v$ where $0<x_{b,v}<q_{b,v}$ (the crucial value of $b$),
  • followed by a (possibly empty) suffix of values $v$ with $x_{b,v}=q_{b,v}>0$ (the tight values of $b$).
Proof Sketch..
The proof relies on a value exchange argument, showing that there is an optimal EA-LP solution where for any two buyers $b^{\prime}$ and $b^{\prime\prime}$, with $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$, and for any two values $v^{\prime\prime}\geq v^{\prime}$, either $x_{b^{\prime},v^{\prime}}=0$ or $x_{b^{\prime\prime},v^{\prime\prime}}=q_{b^{\prime\prime},v^{\prime\prime}}$ (or both – the argument also applies to the special case where $b^{\prime}$ coincides with $b^{\prime\prime}$ and $v^{\prime\prime}>v^{\prime}$, which is relevant to this lemma). Otherwise, we could decrease $x_{b^{\prime},v^{\prime}}$ by $\delta=\min\{x_{b^{\prime},v^{\prime}},q_{b^{\prime\prime},v^{\prime\prime}}-x_{b^{\prime\prime},v^{\prime\prime}}\}$, while increasing $x_{b^{\prime\prime},v^{\prime\prime}}$ by $\delta$. Thus, we increase the objective value of EA-LP, if $v^{\prime\prime}>v^{\prime}$ (or keep it unchanged, if $v^{\prime\prime}=v^{\prime}$), without violating any constraints. The item capacity constraints, in particular, remain satisfied, due to the assumption that $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$. The formal proof of Lemma 5.1 is based on this argument and can be found in Appendix F. ∎
Fractional Allocation Aggregated by Bundle.
To turn the intuition above into an anonymous pricing scheme, we aggregate the allocation of EA-LP by bundle and offer identical menu choices to all buyers demanding the same bundle. We recall that for any bundle $S$, $\mathcal{N}_{S}=\{b:S_{b}=S\}$ denotes the set of buyers interested in $S$. For each value $v\in\mathcal{U}$, we let $x_{S,v}=\sum_{b\in\mathcal{N}_{S}}x_{b,v}$ be the total fractional allocation of $S$ to buyers with realized value $v$.
As before, we call a value $v\in\mathcal{U}$ tight for a bundle $S$ if $x_{S,v}=q_{S,v}>0$, and crucial, if $0<x_{S,v}<q_{S,v}$. For each bundle $S$, we consider the vector $X^{(S)}=(x_{S,v})_{v\in\mathcal{U}}$, whose entries correspond to the optimal fractional allocation of bundle $S$ at different values, ordered in increasing order of $v$. In Appendix F, we prove the following on the structure of $X^{(S)}$.
@@
  • a (possibly empty) prefix of values $v$ where $x_{S,v}=0$,
  • followed by at most one value $v$ with $0<x_{S,v}<q_{S,v}$ (the crucial value of $S$),
  • followed by a (possibly empty) suffix of values $v$ with $x_{S,v}=q_{S,v}$ (the tight values of $S$).
Our pricing menu is built around the important value of a bundle $S$, denoted $w_{S}$, which is the largest value $w$ where $x_{S,w}<q_{S,w}$. If $x_{S,v}=q_{S,v}$ for all $v\in\mathcal{U}$, $S$’s important value is $w_{S}=0$. We note that $w_{S}$ is the crucial value of $S$, if one exists, and $S$’s largest non-tight value (or 0), otherwise.
6 General Framework for Static and Anonymous Bundle Pricing
We next present our general algorithmic framework, which serves as the basis for our static and anonymous bundle pricing mechanisms in the subsequent sections. At a high level, our framework proceeds in the following three steps.
Step 1: Scale Down EA-LP.
We first solve EA-LP, but with the item capacity constraints scaled down by a factor $\gamma>1$ depending on the setting; i.e. $c(e)$ in the first EA-LP constraint becomes $c(e)/\gamma$. We let $\textnormal{{FracOpt}}_{\gamma}$ denote the new fractional optimum and observe that $\textnormal{{FracOpt}}_{\gamma}\geq\textnormal{{FracOpt}}/\gamma$.
Step 2: Construction of the Pricing Menu $M$.
We next create a (possibly randomized) menu $M$ from the solution of EA-LP, using Algorithm 1. Our menu is randomized in the sense that the number of copies of a bundle $S$ at its crucial price may be a random variable (whose value is determined before any buyer arrives); importantly, the menu itself does not contain lotteries. The menu construction follows the structure of EA-LP’s solution described by Lemma 5.2. More specifically, we do not offer any copies of a bundle $S$ at values $v$ with $x_{S,v}=0$. We offer $|\mathcal{N}_{S}|$ (i.e., practically unlimited) copies of $S$ at $S$’s smallest tight value $w_{S}+1$ in line 3 (so, availability of $S$ at price $w_{S}+1$ is only restricted by item capacities). Finally, if $S$’s important value $w_{S}$ is crucial, we offer a limited (and carefully chosen) number of additional copies of $S$ at price $w_{S}$. The number of such copies depends on $x_{S,w_{S}}$ and is the only case where we may use randomization.
Input: Optimal solution $x$ of the $\gamma$-scaled-down EA-LP
for each (desired) bundle $S$ do
   Let $w$ be the important value of $S$ (w.r.t. $x$);
   Offer $|\mathcal{N}_{S}|$ (i.e., practically unlimited) copies of $S$ at price $w+1$;
 if $S$ has a crucial value (which must be $w$) and $x_{S,w}\cdot w>\sum_{v\in\mathcal{U}:v>w}x_{S,v}\cdot v$ then
    if $x_{S,w}>1$ then
         Offer $\lfloor x_{S,w}\rfloor$ additional copies of $S$ at price $w$;
    else
         With probability $\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$, offer one additional copy of $S$ at price $w$;
      end if
   end if
end for
@@ 6 General Framework for Static and Anonymous Bundle Pricing
Step 3: Online Algorithm.
We solve EA-LP and fix a pricing menu $M$, as described in steps 1 and 2 above. The values $v_{b}$ of all buyers $b\in\mathcal{N}$ are realized and an instance $I$ is fixed. Given $M$ and $I$, the almighty adversary determines the arrival order of the buyers in $\mathcal{N}$. Then, the online algorithm operates in a simple posted-pricing setting as follows:
@@
    Buyers arrive sequentially. When a buyer $b$ arrives, they examine the current menu $M$ and choose a subset $J$ (possibly empty) of available menu entries that maximizes their utility.
@@
    If $J\neq\emptyset$, for each chosen menu entry $(S_{j},p_{j},c_{j})\in J$, we decrease its remaining multiplicity $c_{j}$ by $1$. If the residual capacities of all demanded items $e\in\bigcup_{j\in J}S_{j}$ are positive, we decrease all of them by $1$, buyer $b$ is deemed accepted and the algorithm’s social welfare increases by $v_{b}$.
We let $\textnormal{{Alg}}_{M,I}$ denote either the set of buyers accepted by our algorithm or their total social welfare; the intended meaning will be clear from the context. Our goal is to prove that $\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\geq\alpha\cdot\textnormal{{FracOpt}}$ for some $\alpha>0$.
Price Subadditivity and Tie-Breaking.
We always assume that in step (3.1) above, every buyer $b$ selects a single menu entry that corresponds to their requested bundle $S_{b}$. In Appendix B, we show that this assumption is without loss of generality. More precisely, we show that our menu prices are subadditive; hence no buyer $b$ interested in $S_{b}$ has an incentive to select more menu entries whose union includes $S_{b}$. Moreover, we always assume that buyers break ties in favour of acceptance. Thus, if their value is $v$, they select a menu entry at price $v$. In Appendix B, we remove this assumption with a small price perturbation.
6.1 The Unconstrained Algorithm and Its Properties
We now introduce the Unconstrained Algorithm, denoted UAlg, which completely ignores item capacities and is only restricted by the menu entries. Although the menu may induce total demand that exceeds some item capacities, UAlg still accepts all such requests. The exact description of UAlg can be derived from that of the online algorithm, in Step 3 above, by removing the phrase “If the residual capacities $\ldots$ of them by 11” from step (3.2) (since UAlg ignores item capacities).
UAlg is used only for the analysis. It may not be feasible, but it allows us to lower bound the expected social welfare of the online (capacity-feasible) algorithm. We use $\textnormal{{UAlg}}_{M,I}$ to denote either the set of buyers accepted by the unconstrained algorithm or their total social welfare. UAlg is just an analytical tool; it is supposed to operate on the same instance $I$, with the same buyer arrival order and with the same menu $M$ as the actual online algorithm Alg. Therefore, every buyer that belongs to $\textnormal{{Alg}}_{M,I}$ must also belong to $\textnormal{{UAlg}}_{M,I}$.
We start with the first key property of UAlg, proven in Appendix C.